Circular No. 8916

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 C1 (CHEN-GAO) The times of the Jan. 30 and 31 observations published on IAUC 8915

were given in error by the Chinese observers; the correct times, along with more careful measurements of position, are provided below.

2008 UT  α_{2000}  δ_{2000}  Mag.
Jan. 30.58240  22^{h}51^{m}06.^{s}89  +62^{o}15′22′′.0  15.4
31.59073  22 57 05.01  +62 14 39.3  15.7
Feb. 1.66713  23 03 33.84  +62 13 07.1  16.0
2.64737  23 09 33.45  +62 11 19.2  16.6

The following revised parabolic orbital elements are taken from MPEC 2008-C21.

T = 2008 Apr. 16.835 TT  ω = 180.^{o}956
Ω = 307.802 }2000.0
q = 1.26166 AU  i = 61.775

Visual total-magnitude and coma-diameter estimates: Feb. 3.74 UT, 13.3, 0′.8 (M. Lehky, Hradec Kralove, Czech Republic, 0.42-m reflector); 3.76, 12.9, 0′.3 (W. Hasubick, Buchloe, Germany, 44-cm reflector); 3.86, 13.5, 0′.7 (V. Nevski, Vitebsk, Belarus, 0.3-m reflector); 4.89, 12.8, 1′.2 (J. J. Gonzalez, Asturias, Spain, 0.20-m reflector).

EDITORIAL NOTICE Subscribers to the printed IAU Circulars and to the electronic IAUCs

and CBETs are advised that, effective 2008 Mar. 1, paid subscriptions will no longer be an option at the 18-month level. To reduce complications, the maximum allowed subscription term after March 1 will be 12 months. We apologize for the sudden inability in Dec. 2007 to process credit-card payments; a new procedure for processing credit-card payments is being sought, but this could take some months. In the meantime, subscribers are requested to send checks (or postal money orders) payable in U.S. dollars and drawn on U.S. banks; bank “wire” transfers of money are also possible, but are less desirable because they are more complicated for the CBAT to process, and subscribers doing bank transfers must add $20.00 to cover bank fees incurred by the Smithsonian Institution.

2008 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green